# Examining Inequality in Park Quality for Promoting Health Across 35 Global Cities

### v1

@@ 4.6 Quantifying Disparities of Park Scores
One outcome of our study was to quantify disparities of health benefits offered by different parks within a city (Section 2.3, Centralization of Park Quality Within Cities). To measure the disparities in the presence of amenities and facilities associated with health-promoting activities within a city, we propose the following disparity index. The metric essentially quantifies the inequality of the park health scores, as generally, one could expect that good park management would provide for a similar amount of features and facilities in all parks of a city. Since the park health scores could be negative, we can not directly use a standard inequality metric, such as the Gini Index, but had to min-max normalize the park score before computing the Gini Index (Equation 3).
@@
where $X_{Act}$ is the score of activity category ${Act}\in[physical,cultural,ect.]$ of a park p in city c, and
$X^{\prime}=\frac{X-X_{max}}{X_{max}-X_{min}}(min-max\leavevmode\nobreak\ normalization),$
@@ 4.7 Semantic Matching of Flickr Labels and OSM Tags
In our validation (Section 2.2), we used a global dataset of geotagged photos from Flickr, from which we selected all 10,711,513 images that were taken within one of the parks from 35 cities in our study. These images came with user-generated labels and were also partially annotated with computer vision labels from a computer vision algorithm[192, 193]. To obtain semantically equivalent representations of Flickr labels and OSM tags, we employed Sentence-BERT (S-BERT[203]) for text embeddings. We formulated this task as an asymmetric semantic search problem, where the Flickr label was the search term, and the goal was to find the closest matching OSM tag. Given the worldwide reach of our study, the multiple languages present in the user-generated Flickr labels created a challenge in mapping them to the corresponding OSM tags, which were all in English. To address this, we identified the top three languages besides English used in the tags of each city, using the Google MediaPipe[204] Language Detection Model[205]. To ensure that the language detection was accurate and to eliminate named entities, we only used labels where the language detection indicated a confidence of 50% or more. We then translated those tags to English using the respective OPUS machine translation models[206].
A detailed review of the matched pairs revealed that, as anticipated, most pairings were logical based on the text similarity between labels and tags with definitions. However, some minor adjustments were still needed, as some matches were not entirely consistent with the theme of health-promoting activities in parks. For example, the term “outdoor” was initially linked to swimming_pool=outdoor. But as there cannot be a suitable equivalent for “outdoor” on OSM, we removed this pairing and equivalent ones, such as “park,” as all photos were taken in parks. Another instance was the pairing of “water” with substance=water. This match did not reflect the specific use of water features in parks to promote health, so we manually adjusted this to water=river, which better represents bodies of water in parks. Through this review step, we improved the quality of the matched pairs, ensuring they more closely align with the theme of health-promoting park activities. The need for this manual step should not diminish the effectiveness of the semantic search within sentence embeddings. It was merely to eliminate labels that could not meaningfully correspond with an OSM tag and to match a few labels with more domain-relevant tags. This matching process yielded 2,171 label-tag pairs in total. Of these, 1,432 pairs corresponded to an OSM tag with health-promoting benefits, such as “steeplechase” being matched to athletics=steeplechase involving physical health benefits, while 739 pairs, such as “Lamp Post” being matched to man_made=lamp_post did not imply health benefits.
Having assured that the matchings are accurate, we proceeded to profile the parks based on the activities associated with the matched OSM tags, following the same scoring approach as what we used for the OSM park elements and spaces (cf. Equation 2). In our validation, we chose a minimum of 250 images from each park and at least 15 parks in each city. This criterion was established to secure a robust number of images for each park, enhancing the accuracy of our analysis. This approach helped us avoid any potential bias that could have been introduced by individual photographers if a park had only a few images. Figure 15 in the Appendix depicts the linear model of the Flickr labels for London, UK.
@@
The replication package contains tables of the park health scores in the cities: https://github.com/LinusDietz/Health-Promoting-Parks-Replication. The original OpenStreetMap data used for scoring the parks is publicly available and can be best obtained from one of the third-party download servers, for example from https://download.geofabrik.de. The Flickr dataset for the validation can not be shared due to terms of conditions of this dataset.
Code availability
The Python and R code to compute park health scores is available at https://github.com/LinusDietz/Health-Promoting-Parks-Replication.
@@ Appendix A Supplementary Results
A.1 Top Parks by City
The following Table 8 lists the best park of each city by their scores on each activity category.
@@ A.1 Top Parks by City
Amsterdam Sloterpark, 2.2 Sloterpark, 2.38 Park Frankendael, 0.88 Westerpark, 2.18 Amstelpark, 2.29
@@
Auckland Lloyd Elsmore Park, 4.07 Point England Park, 4.52 Paneke / Radonich Park, 2.73 Ambury Regional Park, 3.14 Albert Park, 2.57
@@
Berlin Tempelhofer Feld, 5.3 Landschaftspark Johannisthal/Adlershof, 4.02 Gärten der Welt, 2.88 Tempelhofer Feld, 2.49 Treptower Park, 4.14
@@
Boston Franklin Park, 3.64 Charles River Esplanade, 3.62 Temple Street Park, 1.65 Charles River Esplanade, 3.22 Georges Island, 2.25
@@
Buenos Aires Parque Indoamericano, 4.36 Parque de la Memoria, 3.85 Paseo Arzoumanian, 1.69 Parque de las Ciencias, 2.09 Parque Avellaneda, 3.67
Chicago Lincoln Park, 2.75 Northerly Island, 2.57 Grant Park, 2.16 Lincoln Park, 4 Grant Park, 1.91
@@
Christchurch Canterbury Agricutural Park, 4.56 Bottle Lake Forest Park, 4.13 Hagley Park North, 3.84 Avon River Precinct, 3.5 Hagley Park North, 2.65
Copenhagen Fælledparken, 1.89 Østre Anlæg, 1.55 Husum Bypark, 1.36 Enghaveparken, 1.23 Østre Anlæg, 0.98
@@
Hong Kong {CJK} UTF8min 九龍仔公園 Kowloon Tsai Park, 4.29 {CJK} UTF8min 藝術公園Art Park, 3.92 {CJK} UTF8min 佐敦谷公園 Jordan Valley Park, 2.23 {CJK} UTF8min 迪欣湖活動中心 Inspiration Lake Recreation Centre, 4.79 {CJK} UTF8min灣仔臨時海濱花園 Wan Chai Temporary Promenade, 2.77
Houston Hermann Park, 3.65 Hermann Park, 2.55 Wright-Bembry Park, 1.66 Hermann Park, 2.21 Hermann Park, 2.74
@@
London Old Deer Park, 5.58 Russia Dock Woodland, 4.62 Bushy Park, 3.32 Richmond Park, 3.25 Alexandra Park, 2.37
@@
Madrid Parque Agustín Rodríguez Sahagún, 4.33 Dehesa de la Villa, 6.59 Parque del Retiro, 3.36 Finca Vista Alegre, 4.18 Parque del Retiro, 2.9
@@
Melbourne Albert Park, 4.06 Grant Reserve, 4.19 Fitzroy Gardens, 1.21 Albert Park, 2.45 Carlton Gardens, 2.49
@@
Montreal Parc Jean-Drapeau, 4.73 Parc Angrignon, 5.64 Jardin botanique de Montréal, 2.46 Vieux-Port, 2.68 Parc Jean-Drapeau, 2.93
Moscow Парк Останкино, 3.9 Бирюлёвский дендропарк, 2.58 Выставка достижений народного хозяйства, 2.91 Тимирязевский парк, 4.82 Выставка достижений народного хозяйства, 3.22
@@
New York Pelham Bay Park, 3.52 Prospect Park, 3.23 Bronx Park, 3.7 Brooklyn Bridge Park, 3.76 Fort Tilden, 3.8
Paris Bois de Boulogne, 6.83 Bois de Vincennes, 5.23 Jardin des Plantes, 2.36 Bois de Boulogne, 3.2 Bois de Boulogne, 3.61
@@
Perth Altone Park, 3.94 Kings Park, 4.91 Hyde Park, 2.98 Christ Church Grammar Playing Fields, 2.98 Victoria Gardens, 1.7
@@
Philadelphia East Fairmount Park, 3.49 East Fairmount Park, 3.87 East Fairmount Park, 1.73 Race Street Pier, 2.72 Fort Mifflin, 3.19
@@
Rio De Janeiro Aterro do Flamengo, 5.71 Quinta da Boa Vista, 6.14 Largo da Carioca, 1.02 Campo de Santana, 1.42 Praça Luís de Camões, 2.04
@@
Rome Villa Borghese, 3.89 Villa Glori, 2.85 Parco Agricolo di Casal del Marmo, 2.5 Riserva Naturale dell’Acquafredda, 2.5 Villa Borghese, 2.98
San Diego Balboa Park, 3.35 Mission Bay Park, 2.46 Balboa Park, 1.49 Balboa Park, 3.36 Balboa Park, 2.69
San Francisco Presidio of San Francisco, 2.9 Lake Merced Park, 3.7 Golden Gate Park, 3.11 Presidio of San Francisco, 3.62 Golden Gate Park, 2.99
@@
Seattle Warren G. Magnuson Park, 4.76 Seward Park, 3.7 Washington Park Arboretum, 4.95 Seattle Center, 2.77 Seattle Center, 4.61
Seoul {CJK} UTF8mj올림픽공원, 3.91 {CJK} UTF8mj매봉산공원, 3.45 {CJK} UTF8mj서울숲, 2.48 {CJK} UTF8mj송파나루공원, 3.17 {CJK} UTF8mj경복궁, 4.19
@@
Singapore Changi Business Park, 3.72 Windsor Nature Park, 3.6 Singapore Botanic Gardens, 2.52 The Lawn@Marina Bay, 2.45 Singapore Botanic Gardens, 3.16
St Petersburg парк Героев-Пожарных, 2.82 Парк-дендрарий Ботанического сада Петра Великого, 2.98 Летний сад, 1.68 Приморский парк Победы, 3.73 парк Лесотехнической академии, 4.15
@@
Stockholm Årstafältet, 1.74 Kungsträdgården, 1.6 Sveaplan, 2.45 Karlaplan, 1.26 Humlegården, 1.01
@@
Sydney Centennial Park, 6.9 Centennial Park, 4.16 Sydney Park, 1.68 Lawrence Hargrave Reserve, 2.37 Clarkes Point Reserve, 3.21
Taipeh {CJK} UTF8min 大安森林公園, 6.47 {CJK} UTF8min 關渡自然公園, 6.23 {CJK} UTF8min 士林官邸公園, 2.52 {CJK} UTF8min 天母運動公園, 3.34 {CJK} UTF8min中正紀念公園, 4.14
Tokyo {CJK} UTF8min 若洲海浜公園, 6.14 {CJK} UTF8min上野恩賜公園, 6.77 {CJK} UTF8min北の丸公園, 4.41 {CJK} UTF8min上野恩賜公園, 3.54 {CJK} UTF8min 上野恩賜公園, 7.92
@@
Toronto Centennial Park, 3.26 Sunnybrook Park, 4.54 Highland Creek Ravine, 2.86 Toronto Island Park, 4.32 Don Valley Brick Works Park, 4.02
Vancouver Connaught Park, 1.98 Stanley Park, 3.52 Stanley Park, 1.13 Hastings Park, 2.53 Morton Park, 1.4
Vienna Augarten, 3.63 Draschepark, 3.05 Schlosspark Schönbrunn, 1.45 Schlosspark Schönbrunn, 3.09 Schlosspark Schönbrunn, 2.89
Washington Dc East Potomac Park, 4.28 National Mall, 3.3 National Mall, 3.08 National Mall, 3.98 National Mall, 4.24
A.2 Average Park Offerings by City
Figure 8 subdivides Figure 2 by cities offering a more detailed picture of the average park offering in the cities.
Figure 8: Area of park spaces (horizontal axis) and count of elements (vertical axis) in a fictional average 8-hectare park dedicated to individual activities. Nature-appreciation and physical activity spaces are the most common, with elements dedicated to nature appreciation outnumbering those for social and environmental activities. This plot subdivides Figure 2 by cities. Note that for better readability, the scaling of the axes is not fixed.
@@ A.4 Disparities of Park Scores
Figure 10 provides a visual representation of the data tabulated in Table 6 emphasizing the continent of the cities. This figure supplements the findings in Table 6 The spatial variance contributes to the overall unequal distribution of activity opportunities.
@@ A.5 Validating the Overall Ranking of Parks through an Online Survey
As an additional means for validating the park scores, we conducted a survey in one city, London. In an online questionnaire, we asked London citizens about suitable parks for performing activities. The main set of questions was phrased as: “Can you name several parks suitable for physical activities (e.g., sports)?
@@ Results
Figure 11: The online survey results as box-whisker plots. Standard box plot with Q1, median, Q3; whiskers at 1.5×1.5$\times$1.5 × interquartile range.The values indicated with the x represent the mean values.
@@ Appendix B Supplementary Material
B.1 Determining the Threshold Values for Computing the Linear Models

### v2

@@ 4.6 Quantifying Disparities of Park Scores
One goal of our study was to quantify disparities in health benefits offered by different parks within a city. To measure the disparities in the presence of amenities and facilities associated with health-promoting activities within a city, we propose the following disparity index. The metric essentially quantifies the inequality of the park health scores, as generally, one could expect that good park management would provide for a similar amount of features and facilities in all parks of a city. Since the park health scores could be negative, we can not directly use a standard inequality metric, such as the Gini Index, but had to min-max normalize the park score before computing the Gini Index (Equation 3).
@@
where $X_{Act}$ is the score of activity category ${Act}\in[\text{physical, cultural, etc.}]$ of a park P in city C, and
$X^{\prime}=\frac{X-X_{max}}{X_{max}-X_{min}}(min-max\penalty 10000\ normalization),$
@@ 4.7 Semantic Matching of Flickr Labels and OSM Tags
In our validation, we leveraged a global dataset of 10.7 million geotagged Flickr images taken within parks across 35 cities. Each image came with user-generated tags, partially annotated by computer vision algorithms. To semantically match these Flickr labels to OSM tags, we used text embeddings, treating the task as an asymmetric semantic search problem. To overcome language diversity in the Flickr labels, we detected the top three non-English languages per city and translated the labels into English using machine translation models. To further improve embedding quality, we enriched OSM tags with concise definitions from the OSM mapping guidelines. The embeddings were generated using the all-mpnet-base-v2 S-BERT model, and matching was done based on cosine similarity, with a threshold of $0.7$ to ensure quality. The methodological details are described in Section B.7.
This process yielded $2,171$ Flickr-to-OSM matches, with $1,432$ corresponding to health-promoting features. To assess accuracy, three experts reviewed the $20$ most frequent label-tag matches for London. We aggregated their responses using majority voting. The experts’ annotation agreed with $82\%$ of the matchings, which is highly accurate considering they are based solely on individual tags.
Having assured that the matchings are accurate, we proceeded to profile the parks based on the activities associated with the matched OSM tags, following the same scoring approach as what we used for the OSM park elements and spaces (Equation 1). In our validation, we chose a minimum of $250$ images from each park and at least $15$ parks in each city. This criterion was established to secure a robust number of images for each park, thus ensuring the accuracy of our analysis. This was a mitigation against potential biases that could have been introduced by individual photographers if a park had only a few images.
@@
The replication package contains tables of the park health scores in the cities: https://github.com/LinusDietz/Health-Promoting-Parks-Replication. The original OpenStreetMap data used for scoring the parks is publicly available and can be best obtained from one of the third-party download servers, for example from https://download.geofabrik.de. The Flickr dataset for the validation can not be shared due to the terms of conditions of this dataset.
@@ Appendix A Supplementary Results
A.1 Lexicon of Elements and Spaces for Health-promoting Activities in Parks
We display the top 10 most frequent tags for each activity category in Table 3 and have published all 1441 entries of the lexicon in the replication package.
Activity Category Elements Spaces
A.2 Top Parks by City
The following Table 4 lists the best-equipped park of each city by their scores on each activity category.
@@ A.2 Top Parks by City
Amsterdam Sloterpark Sloterpark Park Frankendael Westerpark Amstelpark
@@
Auckland Lloyd Elsmore Park Point England Park Paneke / Radonich Park Ambury Regional Park Albert Park
@@
Berlin Tempelhofer Feld Landschaftspark Johannisthal/Adlershof Gärten der Welt Tempelhofer Feld Treptower Park
@@
Boston Franklin Park Charles River Esplanade Temple Street Park Charles River Esplanade Georges Island
@@
Buenos Aires Parque Indoamericano Parque de la Memoria Paseo Arzoumanian Parque de las Ciencias Parque Avellaneda
Chicago Lincoln Park Northerly Island Grant Park Lincoln Park Grant Park
@@
Christchurch Canterbury Agricutural Park Bottle Lake Forest Park Hagley Park North Avon River Precinct Hagley Park North
Copenhagen Fælledparken Østre Anlæg Husum Bypark Enghaveparken Østre Anlæg
@@
Hong Kong 九龍仔公園 Kowloon Tsai Park 藝術公園Art Park 佐敦谷公園 Jordan Valley Park 迪欣湖活動中心 Inspiration Lake Recreation Centre 灣仔臨時海濱花園 Wan Chai Temporary Promenade
Houston Hermann Park Hermann Park Wright-Bembry Park Hermann Park Hermann Park
@@
London Old Deer Park Russia Dock Woodland Bushy Park Richmond Park Alexandra Park
@@
Madrid Parque Agustín Rodríguez Sahagún Dehesa de la Villa Parque del Retiro Finca Vista Alegre Parque del Retiro
@@
Melbourne Albert Park Grant Reserve Fitzroy Gardens Albert Park Carlton Gardens
@@
Montreal Parc Jean-Drapeau Parc Angrignon Jardin botanique de Montréal Vieux-Port Parc Jean-Drapeau
Moscow Парк Останкино Бирюлёвский дендропарк Выставка достижений народного хозяйства Тимирязевский парк Выставка достижений народного хозяйства
@@
New York Pelham Bay Park Prospect Park Bronx Park Brooklyn Bridge Park Fort Tilden
Paris Bois de Boulogne Bois de Vincennes Jardin des Plantes Bois de Boulogne Bois de Boulogne
@@
Perth Altone Park Kings Park Hyde Park Christ Church Grammar Playing Fields Victoria Gardens
@@
Philadelphia East Fairmount Park East Fairmount Park East Fairmount Park Race Street Pier Fort Mifflin
@@
Rio De Janeiro Aterro do Flamengo Quinta da Boa Vista Largo da Carioca Campo de Santana Praça Luís de Camões
@@
Rome Villa Borghese Villa Glori Parco Agricolo di Casal del Marmo Riserva Naturale dell’Acquafredda Villa Borghese
San Diego Balboa Park Mission Bay Park Balboa Park Balboa Park Balboa Park
San Francisco Presidio of San Francisco Lake Merced Park Golden Gate Park Presidio of San Francisco Golden Gate Park
@@
Seattle Warren G. Magnuson Park Seward Park Washington Park Arboretum Seattle Center Seattle Center
Seoul 올림픽공원 매봉산공원 서울숲 송파나루공원 경복궁
@@
Singapore Changi Business Park Windsor Nature Park Singapore Botanic Gardens The Lawn@Marina Bay Singapore Botanic Gardens
St Petersburg парк Героев-Пожарных Парк-дендрарий Ботанического сада Петра Великого Летний сад Приморский парк Победы парк Лесотехнической академии
@@
Stockholm Årstafältet Kungsträdgården Sveaplan Karlaplan Humlegården
@@
Sydney Centennial Park Centennial Park Sydney Park Lawrence Hargrave Reserve Clarkes Point Reserve
Taipeh 大安森林公園 關渡自然公園 士林官邸公園 天母運動公園 中正紀念公園
Tokyo 若洲海浜公園 上野恩賜公園 北の丸公園 上野恩賜公園 上野恩賜公園
@@
Toronto Centennial Park Sunnybrook Park Highland Creek Ravine Toronto Island Park Don Valley Brick Works Park
Vancouver Connaught Park Stanley Park Stanley Park Hastings Park Morton Park
Vienna Augarten Draschepark Schlosspark Schönbrunn Schlosspark Schönbrunn Schlosspark Schönbrunn
Washington DC East Potomac Park National Mall National Mall National Mall National Mall
@@ A.3 Geographic Influence on Park Scores
Table 5: Park health scores by quartiles determined using the distance to the city center. Q1 are the inner city parks, and Q4 are the parks that are most distant to the city center. On the left, we show the mean health scores for each quartile; on the right, we show the $p$-values determined using a two-sided t-test for each null hypothesis $H_{0}:Q_{i}=Q_{j}$ stating that the mean scores for two subsequent quartiles $i$ and $j$ are equal. All null hypotheses can be refuted with high significance levels ($p<0.001,\ast\ast\ast$), with the exception of the difference between Q3 and Q4 in the environmental category, where the significance is $p=0.003,\ast\ast$. We confirmed the normal distribution of the scores using $Q$-plots.
Mean Health Scores $p$-values of $H_{0}$
Activity Category Q1 Q2 Q3 Q4 Q1 = Q2 Q2 = Q3 Q3 = Q4
Physical 0.382 0.297 0.160 0.068 0.000 0.000 0.000
Nature-appreciation 0.107 -0.211 -0.497 -0.580 0.000 0.000 0.000
Environmental 0.020 -0.076 -0.156 -0.186 0.000 0.000 0.003
Social 0.289 0.137 -0.004 -0.088 0.000 0.000 0.000
Cultural 0.260 0.096 0.010 -0.022 0.000 0.000 0.001
@@ A.4 Disparities of Park Scores
Figure 6 provides a visual representation of the data tabulated in Table 2 emphasizing the continent of the cities. This figure supplements the findings in Section 2.5.
@@ A.5 Validating the Overall Ranking of Parks through an Online Survey
As an additional means for validating the park scores, we conducted a survey in one city, London, UK. In an online questionnaire, we asked London citizens about suitable parks for performing activities. The main set of questions was phrased as: “Can you name several parks suitable for physical activities (e.g., sports)?
@@ Results
Figure 7: The online survey results as box-whisker plots. Standard box plot with Q1, median, Q3; whiskers at $1.5\times$ interquartile range.The values indicated with the x represent the mean values.
@@ Appendix B Supplementary Material
B.1 Scoping Review to Map Activity Categories to Health Benefits
Results
Activity Category Health Aspect Specific Health Benefit
Physical Cognitive health dementia prevention[69]
General health longevity[70, 5]
Social health various[120, 83], social cohesion[121, 78, 122, 81]
General health lower morbidity[126]
General health lower morbidity[126]
Social health access to healthy produce[161], social cohesion[157, 161]
General health longevity[5]
Cultural Cognitive health dementia prevention[69]
General health various[177]
Physical health increase of physical activity[178]
Well-being quality of life[69, 176, 178]
Mindfulness Mental health stress reduction[139], anxiety reduction[79], depression prevention[82, 79]
Physical health increase of physical activity[114]
Well-being quality of life[179]
A can be seen in Table 7, most research has focused on the benefits of physical activity in parks. Out of 79 studies on the health benefits of exercising, 46 underscored positive outcomes like weight loss[85, 86, 87], cardiovascular improvements[115, 117], metabolic activity[93, 94]. Additionally, these activities demonstrated positive effects on mental health (16 articles), well-being (7 articles), and social health (6 articles).
Finally, cultural and mindfulness activities are relatively under-researched regarding their health benefits in the context of urban parks. Cultural activities often fell outside the scope of our review, which required a connection to urban parks, while more general cultural activities were studied. However, their health benefits are likely underreported given the presence of cultural facilities in parks (e.g., historic monuments or arts venues). Likewise, mindfulness activities and nature-based therapeutic interventions, such as forest therapy[180] provide health benefits, but have not been studied in the context of urban parks, highlighting a potential gap in the scientific literature that warrants future exploration.
In summary, we found that all the different activities we identified have a distinct but overlapping set of health benefits. For example, both physical and nature-appreciation activities help prevent depression, but only physical activities help bone development, and only nature-appreciation brings calmness.
B.2 Determining the Threshold Values for Computing the Linear Models